# Tube-NeRF: Efficient Imitation Learning of Visuomotor Policies from MPC using Tube-Guided Data Augmentation and NeRFs

### 23 Nov 2023 v1

@@ IV-B Tube-guided DA for visuomotor learning
IL objective. We denote the expert output feedback RTMPC in Eq. 6, Eq. 7 and the state observer in Eq. 5 as $\pi_{\theta^{*}}$. The goal is to design an IL and DA strategy to efficiently learn the parameters $\hat{\theta}^{*}$ of the policy Eq. 1, collecting demonstrations from the expert $\pi_{\theta^{*}}$. In IL, this objective is achieved by solving:
$\small\hat{\theta}^{*}\in\text{arg}\min_{\theta}\mathbb{E}_{p(\tau|\pi_{\theta},\mathcal{T})}[J(\theta,\theta^{*}|\tau)],$ (11)
where $J$ is a distance metric, chosen to be the MSE loss:
$\small J(\theta,\theta^{*}|\tau)=\frac{1}{T}\sum_{t=0}^{T-1}\|\pi_{\theta}(\mathbf{o}_{t},\mathbf{X}_{t}^{\text{des}})-\pi_{\theta^{*}}(\mathbf{o}_{t},\mathbf{X}_{t}^{\text{des}})\|_{2}^{2},$ (12)
and $\tau:=\{(\mathbf{o}_{t},\mathbf{u}_{t},\mathbf{X}^{\text{des}}_{t})\!\mid\!\!t\!=\!0,\!\dots,\!T\}$ is a $T+1$ step (observation, action, reference) trajectory sampled from the distribution $p(\tau|\pi_{\theta},\mathcal{T})$. Such a distribution represents all the possible trajectories induced by the student policy $\pi_{\theta}$ in the deployment environment $\mathcal{T}$. As observed in [8, 9], the presence of uncertainties in $\mathcal{T}$ makes IL challenging, as demonstrations are usually collected in a training domain ($\mathcal{S}$) under a different set of uncertainties ($\mathbb{W}_{\mathcal{S}}\subseteq\mathbb{W}_{\mathcal{T}}$, $\mathbb{V}_{\mathcal{S}}\subseteq\mathbb{V}_{\mathcal{T}}$) resulting in a different distribution of training data.
@@ IV-B2 Extra states and actions for synthetic data generation
For every timestep $t$ in $\bar{\tau}$, we generate $N_{\text{synthetic},t}>0$ (details on how $N_{\text{synthetic},t}$ is computed are provided in Section IV-B4) extra (state, action) pairs $(\mathbf{x}_{t,j}^{+},\mathbf{u}_{t,j}^{+})$, with $j=1,\dots,N_{\text{synthetic},t}$ by sampling extra states from the tube $\mathbf{x}_{t,j}^{+}\in\bar{\mathbf{x}}_{t}^{*}\oplus\mathbb{Z}$, and computing the corresponding control action $\mathbf{u}_{t,j}^{+}$ using the ancillary controller Eq. 7:
@@ IV-B3 Synthetic observations generation
To generate the necessary data $\mathbf{o}_{t,j}^{+}=(\mathcal{I}_{t,j}^{+},\mathbf{o}_{\text{other},{t,j}}^{+})$ input for the sensorimotor policy Eq. 1 from the selected states $\mathbf{x}_{j,t}^{+}$, we employ observation models Eq. 4 available for the expert. In the context of learning a visuomotor policy, we generate synthetic camera images $\mathcal{I}_{t,j}^{+}$ using an inverse pose estimator $\hat{g}^{-1}_{\text{\text{cam}}}$, mapping camera poses $\mathbf{T}_{IC}$ to images $\mathcal{I}$ via $\small\mathcal{I}_{t,j}^{+}=\hat{g}_{\text{\text{cam}}}^{-1}(\mathbf{T}_{{IC}_{t,j}}^{+}),$ where $\mathbf{T}_{IC}$ denotes a homogeneous transformation matrix from a world (inertial) frame $I$ to a camera frame $C$. $\hat{g}^{-1}_{\text{\text{cam}}}$ is obtained by generating a NeRF of the environment (discussed in more details in Section V-1) from the images $\mathcal{I}_{0},\dots,\mathcal{I}_{T}$ in the collected demonstration $\bar{\tau}$, and by estimating the intrinsic/extrinsic of the camera onboard the robot. The camera poses $\textbf{T}_{{IC}_{t,j}}^{+}$ are obtained from the sampled states $\mathbf{x}_{t,j}^{+}$, which includes the robot’s position and orientation. These are computed as $\textbf{T}_{{IC}_{t,j}}^{+}=\textbf{T}_{IB}(\mathbf{x}_{t,j}^{+})\hat{\textbf{T}}_{{BC}_{t,j}}$, where $\textbf{T}_{IB}$ is the transformation from the robot’s body frame $B$ to the reference frame $I$, and $\hat{\textbf{T}}_{{BC}_{t,j}}$ is the perturbed transformation between $B$ and the camera frame $C$. Perturbations in $\hat{\textbf{T}}_{{BC}_{t,j}}$, applied to the camera’s nominal extrinsic $\textbf{T}_{BC}$, accommodate uncertainties and errors in the camera extrinsics. Last, the full observations $\mathbf{o}_{t,j}^{+}$ are obtained by computing $\mathbf{o}_{\text{other},t,j}^{+}$, using Eq. 4 and a selection matrix $\mathbf{S}$:
@@ V APPLICATION TO VISION-BASED FLIGHT
Task. We apply our framework to learn to track a figure-eight trajectory (lemniscate, with velocity up to $3.15$ m/s lasting $30$s), denoted T1. Robot model. The expert uses the hover-linearized model of a multirotor [34], with state $\mathbf{x}=[_{I}\bm{p}^{T},_{I}\bm{v}^{T},_{\bar{I}}\varphi,_{\bar{I}}\vartheta]^{T},$ (position $\bm{p}\in\mathfrak{R}^{3}$, velocity $\bm{v}\in\mathfrak{R}^{3}$, roll $\varphi\in\mathfrak{R}$, pitch $\vartheta\in\mathfrak{R}$, $n_{x}\!=\!8$), $I$ is an inertial reference frame, while ${\bar{I}}$ a yaw-fixed frame [34]. The control input $\mathbf{u}_{t}$ ($n_{u}\!=\!3$) is desired roll, pitch, and thrust, and it is executed by a cascaded attitude controller. Measurements. Our multirotor is equipped with a fisheye monocular camera, tilted $45\deg$ downwards, that generates images $\mathcal{I}_{t}$ (size $128\times 96$ pixels). In addition, we assume available onboard noisy altitude ${}_{I}p^{m}_{z}\in\mathfrak{R}$, velocity ${}_{I}\bm{v}^{m}\in\mathfrak{R}^{3}$ and roll ${}_{I}\varphi^{m}$, pitch ${}_{I}\vartheta^{m}$ and yaw measurements. This is a common setup in aerial robotics, where noisy altitude and velocity can be obtained, for example, via optical flow and a downward-facing lidar, while roll, pitch, and yaw can be computed from an IMU with a magnetometer, using a complementary filter [35]. Student policy. The student policy Eq. 1, whose architecture is shown in Figure 3, takes as input an image $\mathcal{I}_{t}$ from the onboard camera, the reference trajectory $\mathbf{X}^{\text{des}}_{t}$, and $\mathbf{o}_{\text{other}}\coloneqq[_{I}p^{m}_{z},{{}_{I}\bm{v}^{m}}^{T},_{\bar{I}}\varphi^{m},_{\bar{I}}\vartheta^{m}]^{T},$ and it outputs $\mathbf{u}_{t}$. A Squeezenet [36] is used to map $\mathcal{I}_{t}$ into a lower-dimensional feature space; it was selected for its performance at a low computational cost. To promote learning of internal features relevant to estimating the robot’s state, the output of the policy is augmented to predict the current state $\hat{\mathbf{x}}$ (or $\mathbf{x}^{+}$ for the augmented data), modifying Eq. 12 accordingly. This output was not used at deployment time, but it was found to improve the performance. Output Feedback RTMPC and observer. The expert uses the defined robot model for predictions, discretized with $T_{s}=0.1$ s, and horizon $N=30$, ($3.0$ s). $\mathbb{X}$ encodes safety and position limits, while $\mathbb{U}$ captures the maximum actuation capabilities. Process uncertainty in $\mathbb{W}_{\mathcal{T}}$ is assumed to be a bounded external force disturbance with magnitude up to $20\%$ of the weight of the robot and random direction, close to the physical limits of the platform. The state estimator Eq. 5 is designed by using as measurement model in Eq. 4: $\bar{\mathbf{o}}_{t}=\mathbf{x}_{t}+\mathbf{v}_{t}$ where we assume $\mathbf{v}_{t}\!\sim\!\mathcal{N}(\mathbf{0}_{n_{x}},[\bm{\sigma}^{T}_{\text{cam}},\bm{\sigma}^{T}_{\text{other}}]^{T})$. We therefore set $\mathbb{V}_{\mathcal{T}}=\{\mathbf{v}_{t}\;|\;\|\mathbf{v}\|_{\infty}\leq 3\left[\bm{\sigma}_{\text{cam}}^{T},\bm{\sigma}_{\text{other}}^{T}\right]^{T}\}$, with $3\bm{\sigma}_{\text{cam}}=[0.6,0.6]^{T}$ (units in $m$) and $3\bm{\sigma}_{\text{other}}=[0.4,0.2,0.2,0.2,0.05,0.05]^{T}$ (units in $m$ for altitude, $m/s$ for velocity, and $rad$ for tilt). These conservative but realistic parameters are based on prior knowledge of the worst-case performance of vision-based estimators in our relatively feature-poor flight space. The observer gain matrix $\mathbf{L}$ is computed by assuming fast state estimation dynamics, (poles of $\mathbf{A-LC}$ at $30.0$ rad/s). For simplicity, the RPI set $\mathbb{S}$ is obtained from $\mathbb{W}_{\mathcal{T}}$ and $\mathbb{V}_{\mathcal{T}}$ via Monte Carlo simulations of Eq. 8, uniformly sampling instances of the uncertainties, and by computing an outer axis-aligned bounding box of the trajectories of $\bm{\xi}$.
@@ V-1 Procedure to generate the NeRF of the environment
(i) Dataset:A NeRF of the environment, the MIT-Highbay, is generated from about $100$ images collected during a single real-world demonstration of the figure-eight trajectory (T1) intended for learning, utilizing full-resolution images ($640\times 480$ pixels) from the fisheye camera onboard the Qualcomm Snapdragon Flight Pro board of our UAV. (ii) Extrinsic/Intrinsic:The extrinsic and intrinsic parameters of the camera are estimated from the dataset using structure-from-motion (COLMAP [37]). We chose the RADTAN camera model, as this was found to provide better reconstructions than higher-order fisheye models. (iii) Frame Alignment:The scale and homogeneous transformation aligning the reference frame used by the COLMAP with the reference frame used by the robot’s state estimator are determined via the trajectory alignment tool EVO [38]. This enables the integration of the NeRF as an image rendering tool in a simulation/DA framework. (iv) NeRF Training:Instant-NGP [27] is utilized to train the NeRF. The scaling of the Axis-Aligned Bounding Box used by the Instant-NGP is manually adjusted to ensure that the reconstruction is photorealistic in the largest possible volume. (v) Images rendering for DA:Novel images are rendered using the same camera intrinsics identified by COLMAP. The camera extrinsics, mapping from the robot’s IMU to the optical surface, are determined via Kalibr [39], using an ad-hoc dataset. An example of an image from the NeRF is shown in Figure 1.
@@ VI-A Evaluation in simulation
Figure 4: Episode length (number of timestep before a state constraint violation, for a maximum of $300$ timesteps) as a function of the number of demonstrations collected from the expert, and as a function of the training time (the time required to collect such demonstrations in simulation, and to train the policy). Two deployment domains are considered, one with sensing noise and one with additional process noise (wind-like disturbances). The results highlight that policies trained with Tube-NeRF (TN) archive full episode length after a single demonstration, and require less than half of the training time than the best-performing baselines (DR-based methods). Reducing the number of samples extracted from the tube (e.g., TN-50) improves the training time. Note that the lines of Tube-NeRF-based approaches vs the number of demonstrations overlap. Shaded areas are $95\%$ confidence intervals. Note that to focus our study on the effects of process uncertainties and sensing noise, we do not apply visual changes to the environment, nor the robustification to visual changes (Section IV-B5). Evaluations across $10$ seeds.
@@
TABLE III: Position RMS Tracking Errors on the ${x,y,z}$ axis across multiple runs when following a figure-eight (lemniscate) trajectory (T1) and a circular trajectory (T2). These results highlight the overall low tracking errors of the sensorimotor policy, comparable to the ones of the expert whose position is estimated via a motion capture system. We apply wind disturbances with a leaf blower (With wind), and extra sensing noise in the altitude, velocity and orientation part input of the policy (High noise). Unlike in T1, the policy trained for T2 has been obtained without collecting corresponding images from a real-world demonstration, and therefore exclusively relies on the NeRF-based data augmentation during certain parts of the trajectory (see Figure 7). All experiments repeated $3$ times, except T2 with wind low noise (two times), and slung-load (once).

### 26 Feb 2024 v2

@@ IV-B Tube-guided DA for visuomotor learning
IL objective. We denote the expert output feedback RTMPC in Eq. 6, Eq. 7 and the state observer in Eq. 5 as $\pi_{\theta^{*}}$. The goal is to design an IL and DA strategy to efficiently learn the parameters $\hat{\theta}^{*}$ of the policy Eq. 1, collecting demonstrations from the expert $\pi_{\theta^{*}}$. In IL, this objective consists in minimizing the MSE loss:
$\small\hat{\theta}^{*}\!\in\!\text{arg}\!\min_{\theta}\!\mathbb{E}_{p(\tau|\pi_{\theta},\mathcal{T})}\!\left[\frac{1}{T}\sum_{t=0}^{T-1}\|\pi_{\theta}(\mathbf{o}_{t},\mathbf{X}_{t}^{\text{des}})\!\!-\!\!\pi_{\theta^{*}}(\mathbf{o}_{t},\mathbf{X}_{t}^{\text{des}})\|_{2}^{2}\right]$
where $\tau:=\{(\mathbf{o}_{t},\mathbf{u}_{t},\mathbf{X}^{\text{des}}_{t})\!\mid\!\!t\!=\!0,\!\dots,\!T\}$ is a $T+1$ step (observation, action, reference) trajectory sampled from the distribution $p(\tau|\pi_{\theta},\mathcal{T})$. Such a distribution represents all the possible trajectories induced by the student policy $\pi_{\theta}$ in the deployment environment $\mathcal{T}$. As observed in [7, 8], the presence of uncertainties in $\mathcal{T}$ makes IL challenging, as demonstrations are usually collected in a training domain ($\mathcal{S}$) under a different set of uncertainties ($\mathbb{W}_{\mathcal{S}}\subseteq\mathbb{W}_{\mathcal{T}}$, $\mathbb{V}_{\mathcal{S}}\subseteq\mathbb{V}_{\mathcal{T}}$) resulting in a different distribution of training data.
@@ IV-B2 Extra states and actions for synthetic data generation
For every timestep $t$ in $\bar{\tau}$, we generate $N_{\text{synthetic},t}>0$ (details on how $N_{\text{synthetic},t}$ is computed are provided in Section IV-B4) extra (state, action) pairs $(\mathbf{x}_{t,j}^{+},\mathbf{u}_{t,j}^{+})$, with $j=1,\dots,N_{\text{synthetic},t}$ by sampling extra states from the tube $\mathbf{x}_{t,j}^{+}\in\bar{\mathbf{x}}_{t}^{*}\oplus\mathbb{Z}$, and computing the corresponding control action $\mathbf{u}_{t,j}^{+}$ using Eq. 7:
@@ IV-B3 Synthetic observations generation
To generate the necessary data $\mathbf{o}_{t,j}^{+}=(\mathcal{I}_{t,j}^{+},\mathbf{o}_{\text{other},{t,j}}^{+})$ input for the sensorimotor policy Eq. 1 from the selected states $\mathbf{x}_{j,t}^{+}$, we employ observation models Eq. 4 available for the expert. In the context of learning a visuomotor policy, we generate synthetic camera images $\mathcal{I}_{t,j}^{+}$ using an inverse pose estimator $\hat{g}^{-1}_{\text{\text{cam}}}$, mapping camera poses $\mathbf{T}_{IC}$ to images $\mathcal{I}$ via $\small\mathcal{I}_{t,j}^{+}=\hat{g}_{\text{\text{cam}}}^{-1}(\mathbf{T}_{{IC}_{t,j}}^{+}),$ where $\mathbf{T}_{IC}$ denotes a homogeneous transformation matrix from a world (inertial) frame $I$ to a camera frame $C$. $\hat{g}^{-1}_{\text{\text{cam}}}$ is obtained by generating a NeRF of the environment (discussed in more details in Section V-1) from the images $\mathcal{I}_{0},\dots,\mathcal{I}_{T}$ in the collected demonstration $\bar{\tau}$, and by estimating the intrinsic/extrinsic of the camera onboard the robot. The camera poses $\textbf{T}_{{IC}_{t,j}}^{+}$ are obtained from the sampled states $\mathbf{x}_{t,j}^{+}$, which includes the robot’s position and orientation. These are computed as $\textbf{T}_{{IC}_{t,j}}^{+}=\textbf{T}_{IB}(\mathbf{x}_{t,j}^{+})\hat{\textbf{T}}_{{BC}_{t,j}}$, where $\textbf{T}_{IB}$ is the transformation from the robot’s body frame $B$ to the reference frame $I$, and $\hat{\textbf{T}}_{{BC}_{t,j}}$ are perturbed transformation of the nominal camera extrinsic, where perturbations are introduced to accommodate uncertainties and errors in the extrinsics. Last, the full observations $\mathbf{o}_{t,j}^{+}$ are obtained by computing $\mathbf{o}_{\text{other},t,j}^{+}$, using Eq. 4 and a selection matrix $\mathbf{S}$:
@@ V APPLICATION TO VISION-BASED FLIGHT
Task. We apply our framework to learn to track a figure-eight trajectory (lemniscate, with velocity up to $3.15$ m/s lasting $30$s), denoted T1. Robot model. The expert uses the hover-linearized model of a multirotor [30], with state $\mathbf{x}=[_{I}\bm{p}^{T},_{I}\bm{v}^{T},_{\bar{I}}\varphi,_{\bar{I}}\vartheta]^{T},$ (position $\bm{p}\in\mathfrak{R}^{3}$, velocity $\bm{v}\in\mathfrak{R}^{3}$, roll $\varphi\in\mathfrak{R}$, pitch $\vartheta\in\mathfrak{R}$, $n_{x}\!=\!8$), $I$ is an inertial reference frame, while ${\bar{I}}$ a yaw-fixed frame [30]. The control input $\mathbf{u}_{t}$ ($n_{u}\!=\!3$) is desired roll, pitch, and thrust, and it is executed by a cascaded attitude controller. Measurements. Our multirotor is equipped with a fisheye monocular camera, tilted $45\deg$ downwards, that generates images $\mathcal{I}_{t}$ (size $128\times 96$ pixels). In addition, we assume available onboard noisy altitude ${}_{I}p^{m}_{z}\in\mathfrak{R}$, velocity ${}_{I}\bm{v}^{m}\in\mathfrak{R}^{3}$ and roll ${}_{I}\varphi^{m}$, pitch ${}_{I}\vartheta^{m}$ and yaw measurements. This is a common setup in aerial robotics, where noisy altitude and velocity can be obtained, for example, via optical flow and a downward-facing lidar, while roll, pitch, and yaw can be computed from an IMU with a magnetometer, using a complementary filter [31]. Student policy. The student policy Eq. 1, whose architecture is shown in Figure 3, takes as input an image $\mathcal{I}_{t}$ from the onboard camera, the reference trajectory $\mathbf{X}^{\text{des}}_{t}$, and $\mathbf{o}_{\text{other}}\coloneqq[_{I}p^{m}_{z},{{}_{I}\bm{v}^{m}}^{T},_{\bar{I}}\varphi^{m},_{\bar{I}}\vartheta^{m}]^{T},$ and it outputs $\mathbf{u}_{t}$. A Squeezenet [32] is used to map $\mathcal{I}_{t}$ into a lower-dimensional feature space; it was selected for its performance at a low computational cost. To promote learning of internal features relevant to estimating the robot’s state, the output of the policy is augmented to predict the current state $\hat{\mathbf{x}}$ (or $\mathbf{x}^{+}$ for the augmented data), modifying the training loss accordingly. This output was not used at deployment time, but it was found to improve the performance. Output Feedback RTMPC and observer. The expert uses the defined robot model for predictions, discretized with $T_{s}=0.1$ s, and horizon $N=30$, ($3.0$ s). $\mathbb{X}$ encodes safety and position limits, while $\mathbb{U}$ captures the maximum actuation capabilities. Process uncertainty in $\mathbb{W}_{\mathcal{T}}$ is assumed to be a bounded external force disturbance with magnitude between to $10-19\%$ of the weight of the robot and random direction, close to the physical limits of the platform, and the tube of the expert is computed assuming $\bar{\mathbb{W}}_{\mathcal{T}}$ equal to $20\%$ of the weight of the robot. The state estimator Eq. 5 is designed by using as measurement model in Eq. 4: $\bar{\mathbf{o}}_{t}=\mathbf{x}_{t}+\mathbf{v}_{t}$ where we assume $\mathbf{v}_{t}\!\sim\!\mathcal{N}(\mathbf{0}_{n_{x}},[\bm{\sigma}^{T}_{\text{cam}},\bm{\sigma}^{T}_{\text{other}}]^{T})$. We therefore set $\mathbb{V}_{\mathcal{T}}=\{\mathbf{v}_{t}\;|\;\|\mathbf{v}\|_{\infty}\leq 3\left[\bm{\sigma}_{\text{cam}}^{T},\bm{\sigma}_{\text{other}}^{T}\right]^{T}\}$, with $3\bm{\sigma}_{\text{cam}}=[0.6,0.6]^{T}$ (units in $m$) and $3\bm{\sigma}_{\text{other}}=[0.4,0.2,0.2,0.2,0.05,0.05]^{T}$ (units in $m$ for altitude, $m/s$ for velocity, and $rad$ for tilt). These conservative but realistic parameters are based on prior knowledge of the worst-case performance of vision-based estimators in our relatively feature-poor flight space. The observer gain matrix $\mathbf{L}$ is computed by assuming fast state estimation dynamics, (poles of $\mathbf{A-LC}$ at $30.0$ rad/s).
@@ V-1 Procedure to generate the NeRF of the environment
(i) Dataset:A NeRF of the environment, the MIT-Highbay, is generated from about $100$ images collected during a single real-world demonstration of the figure-eight trajectory (T1) intended for learning, utilizing full-resolution images ($640\times 480$ pixels) from the fisheye camera onboard the Qualcomm Snapdragon Flight Pro board of our UAV. (ii) Extrinsic/Intrinsic:The extrinsic and intrinsic parameters of the camera are estimated from the dataset using structure-from-motion (COLMAP [33], RADTAN camera model). (iii) Frame Alignment:The scale and homogeneous transformation aligning the reference frame used by the COLMAP with the reference frame used by the robot’s state estimator are determined via the trajectory alignment tool EVO [34]. This enables the integration of the NeRF as an image rendering tool in a simulation/DA framework. (iv) NeRF Training:Instant-NGP [24] is utilized to train the NeRF. The scaling of the Axis-Aligned Bounding Box used by the Instant-NGP is manually adjusted to ensure that the reconstruction is photorealistic in the largest possible volume. (v) Images rendering for DA:Novel images are rendered using the same camera intrinsics identified by COLMAP. The camera extrinsics, mapping from the robot’s IMU to the optical surface, are determined via Kalibr [35], using an ad-hoc dataset. An example of an image from the NeRF is shown in Figure 1.
@@ VI-A Evaluation in simulation
[…]
Figure 4: Episode length (timestep before a state constraint violation, up to $300$) vs. number of demonstrations collected from the expert, and vs the training time (the time required to collect such demonstrations in simulation, and to train the policy). This shows that policies trained with Tube-NeRF (TN) archive full episode length after a single demonstration, and require less than half of the training time than the best-performing baselines (DR-based methods). Note that the lines of Tube-NeRF-based approaches vs the number of demonstrations overlap. Shaded areas are $95\%$ confidence intervals. Note that to focus our study on the effects of process uncertainties and sensing noise, we do not apply visual changes to the environment, nor the robustification to visual changes (Section IV-B5). Evaluations across $10$ seeds, $10$ times per seed.
@@
Evaluation details. For every method, we: (i) collect $K$ new demonstrations ($K=1$ for Tube-NeRF, $K=10$ otherwise) via the output feedback RTMPC expert and the state estimator; (ii) update [^2] Policies are trained for $50$ epochs with the ADAM optimizer, learning rate $0.001$, batch size $32$, and terminating training if the loss does not decrease within $7$ epochs. Tube-NeRF uses only the newly collected demonstrations and the corresponding augmented data to update the previously trained policy. a student policy using all the demonstrations collected so far; (iii) evaluate the obtained policy in the considered target environments, for $10$ times each, starting from different initial states; (iv) repeat from (i). Note that in our comparison the environment steps at its highest possible rate (simulation time is faster than wall-clock time), providing an advantage in terms of data collection time to those methods that require collecting a large number of demonstrations (our baselines). Results. Figure 4 highlights that all Tube-NeRF methods, combined with either DAgger or BC, can achieve complete robustness (full episode length) under combined sensing and process uncertainties after a single demonstration. The baseline approaches require $20$-$30$ demonstrations to achieve a full episode length in the environment without wind, and the best-performing baselines (methods with DR) require about $80$ demonstrations to achieve their top episode length in the more challenging environment. Similarly, Tube-NeRF requires less than half training time than DR-based methods to achieve higher robustness in this more challenging environment, and reducing the number of samples (e.g, Tube-NeRF-$50$) can further improve the training time. The time to generate the NeRF, not shown in Fig. 4, was approximately $5$ minutes ($20000$ epochs on an RTX 3090 GPU). Even accounting for this time, TN is significantly faster than collecting real-world demonstrations (the $80$ demonstrations required by DR correspond to $40$ minutes of real-world data, followed by the time to train the policy). In addition, if the NeRF is combined with a simulation of the dynamics of the robot (creating a photo-realistic simulator), our DA strategy still provides benefits in terms of performance and training time.. Table II additionally highlights the small gap of Tube-NeRF policies from the expert in terms of tracking performance (expert gap), and shows that increasing the number of samples (e.g., Tube-NeRF-$100$) benefits performance.
VI-B Flight experiments
TABLE III: Position RMS Tracking Errors when following a figure-eight (lemniscate) trajectory (T1) and a circular trajectory (T2). This highlights the low errors of the sensorimotor policy, comparable to the ones of the expert whose position is estimated via a motion capture system. We apply wind disturbances with a leaf blower (With wind), and extra sensing noise in the altitude, velocity and orientation input of the policy (High noise). Note that T2 as been obtained without collecting corresponding images from a real-world demonstration, and heavily relies on NeRF data (see Figure 7). Experiments repeated $3$ times, except T2 with wind low noise ($2$ times), and slung-load (once).